OpenStudy (anonymous):

Find the 50th term of the sequence 5, -2, -9, -16, ...

OpenStudy (anonymous):

We're subtracting 7 each time. You can see that, right? To get to the 50th term, we have to subtract 7 49 times. 5 - 49*7 = 5 - 343 = -338

OpenStudy (anonymous):

Thank you :)